OpenStudy (anonymous):

A tea bag is shaped like a regular square pyramid. Each leg of the base is 4 cm, and the slant height is 5 cm. What are the lateral area & surface area of the tea bag? a. 40 cm^2 ; 56 cm^2 b. 40 cm^2 ; 52.6 cm^2 c. 36.6 cm^2 ; 52.6 cm^2 d. 36.7 cm^2 ; 56 cm^2 ive already figured out the lateral area its 40 so it a or b

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

What's the area of the base of the pyramid?

OpenStudy (anonymous):

4^2 so 16?

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

good 

jimthompson5910 (jim_thompson5910):

So total surface area is (Area of base) + (Lateral surface area) = 16+40 = 56 square cm

OpenStudy (anonymous):

ok so a thank you!

jimthompson5910 (jim_thompson5910):

correct, yw